Thursday, 27 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

Marcelo Viana, IMPA, Rio de Janeiro, Brazil

Dynamics: A Probabilistic and Geometric Perspective (I: p.25)

Chair: Jacob Palis, IMPA, Rio de Janeiro, Brazil

10:15 – 11:15  Auditorium Maximum H 105

Stéphane Mallat, École Polytechnique, CMAP, Palaiseau, France and Courant Institute, New York University, New York, USA Applied Mathematics Meet Signal Processing (I: p.13)

Chair: David Mumford, Brown University, Providence, USA

11:45 – 12:45  Auditorium Maximum H 105

Peter Sarnak, Princeton University, Princeton, USA

Zeta and L-Functions (I: p.19)

Chair: John Coates, University of Cambridge, Cambridge, UK

14:00 – 15:00  Auditorium Maximum H 105

Persi Diaconis, Mathematics and ORIE, Cornell University, Ithaca, USA and Departments of Mathematics and Statistics, Stanford University, Stanford, USA From Shuffling Cards to Walking Around the Building (I: p.6)

Chair: David Siegmund, Stanford University, Stanford, USA

Closing Ceremony

15:00 – 16:00  Auditorium Maximum H 105

Program to be announced

Other Events

\underline{17:30 Urania}

Klaus Becker, Klieve, Germany Von der Steinbildhauerei zur Mathematik: Eine Annäherunng an die einbeschriebenen Kugeln regelmäßiger Polyeder (in German)

\underline{18:00 Urania}

Movie Qubba for Al–Kashi / Arabesques and Geometry / Animato

\underline{19:30 Urania}

Heinz-Otto Peitgen, Bremen, Germany Mathematik und Medizin — Computerunterstützung in der Leberchirurgie und Früherkennung von Brustkrebs (in German)

\underline{20:00 Urania} VideoMath Festival

The “VideoMath Reel”, the 24 best contributions to the VideoMath competition